OpenStudy (anonymous):

For the ellipse, find the vertices and foci. 4x^2 + 9y^2 = 1 I mainly need help on finding the vertices so you can ignore the foci part unless you want to answer it. I just need the first 2 - 3 steps for the vertices because I am unsure of how to approach this.

OpenStudy (aric200):

For starters you need to simplify

OpenStudy (anonymous):

How would I simplify? Do I need to factor?

OpenStudy (aric200):

Nevermind i read it wrong hold on

OpenStudy (anonymous):

rewrite the equation in the standard form

OpenStudy (anonymous):

do you know what the standard form is?

OpenStudy (anonymous):

Is it the quadratic standard form or the standard form used when dealing with y=mx+b?

OpenStudy (anonymous):

no, this is the equation of an ellipse we're talking about. x^2/a^2 + y^2/b^2 = 1

OpenStudy (anonymous):

Oh okay. But how would I get a and b into a fraction? Usually I take the last term or number (which in this case is 1) and turn it into a fraction and multiply it to both sides of the equation to get standard form however the 1 in this equation threw me off.

OpenStudy (anonymous):

4x^2 + 9y^2 = 1 is equivalent to x^2/(1/4) + y^2/(1/9) = 1

OpenStudy (anonymous):

what can you do from here?

OpenStudy (anonymous):

thank you very much! I understand it better now!

OpenStudy (anonymous):

ok, no problem